OpenStudy (jolo_yolo):

In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x) 2. Use the table to determine at approximately which point the number of cells will be the same for each organism. 3. Graph the system of equations, and show the point of intersection. 4. Explain what the points graphed for each line represent. 5. Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph from part c

OpenStudy (aripotta):

for number 2, set up the equations so that they're equal to each other, and solve for x 100 + 23x = 90(1.2x)

OpenStudy (jolo_yolo):

to solve for x would i divide both sides by 23x?

OpenStudy (jolo_yolo):

@AriPotta

OpenStudy (aripotta):

no, you'd first multiply 90 and 1.2x

OpenStudy (jolo_yolo):

108x -- 100 + 23x = 108x right?

OpenStudy (aripotta):

right. and now you wanna get rid of the 23x on the left, how should you do that?

OpenStudy (jolo_yolo):

division

OpenStudy (aripotta):

no, we're not dividing yet because we have 2 x's in the equation and we don't want that. so because 23x is being added, we're going to subtract 23x from both sides 100 = 85x

OpenStudy (aripotta):

now we can divide ;)

OpenStudy (jolo_yolo):

oh okay sorry. so now divide both sides by 85x?

OpenStudy (aripotta):

well, divide both sides by 85

OpenStudy (jolo_yolo):

wait, sorry, what does the equation look like now?

OpenStudy (aripotta):

100 = 85x

OpenStudy (jolo_yolo):

okay, just making sure. 1.17 = x?

OpenStudy (aripotta):

i'd round up to 1.18, but yes :)

OpenStudy (jolo_yolo):

alright so that's part 2, right

OpenStudy (aripotta):

yes that was 2

OpenStudy (jolo_yolo):

thank you, can you please help me with the rest? im absolutely stuck

OpenStudy (aripotta):

could you open a new question? i don't like scrolling up and down so long

OpenStudy (jolo_yolo):

yea sure

OpenStudy (aripotta):

alright thanks :)